# Maritime Cybersecurity: A Comprehensive Review

### v2

@@ 5.3. Device heterogeneity
Maritime devices operate in a wide range of environments, ranging from ports and coastal areas and vast expanses of open seas. The diversity of maritime environments requires a corresponding diversity in device types. While this heterogeneity presents challenges for manufacturers and operators alike – including issues related to standardization and interoperability – it also ensures that appropriate technologies can be effectively deployed across various maritime scenarios. The maritime industry is facing an increasing demand for improved operational visibility at sea (Wang et al., 2020). Enhanced operational visibility at sea contributes to overall safety by providing early detection of anomalies or hazards that could pose a threat to crew members or the environment. Maritime IoT addresses this requirement by making it possible to monitor vital indicators like fuel usage, machinery performance, and general vessel operations in real time. As a result, maritime stakeholders are able to make more informed decisions, optimize resource utilization, and improve overall efficiency and safety. However, the implementation of maritime IoT presents complexities. The system must accommodate a wide range of machine-type communication devices, from low-cost units with limited functionality to high-end devices offering advanced features (Wang et al., 2020). Low-cost devices such as sensors and buoys often operate under power and energy constraints, posing challenges in selecting appropriate communication technologies and protocols that prioritize energy efficiency and longevity over high data rates or complex functionalities (Wang et al., 2020). To address these challenges posed by device heterogeneity, edge computing and cloud-based solutions have been proposed (Dhivvya et al., 2017). While cloud-based solutions and edge computing offer valuable tools for managing device heterogeneity in maritime IoT, cloud solutions require reliable internet connections, which may not always be available in open seas where face connectivity issues (Kim et al., 2019). In addition, implementing both cloud and edge solutions adds another layer of complexity to an already complex system of heterogeneous devices.
5.4. Rapid technology evolution v.s. Slow-paced shipping industry
The slow pace of technological adoption in the global shipping fleet is evident in the statistics: only 0.3% of operating vessels worldwide have implemented alternative energy solutions, with this figure rising slightly to 6.05% for ships currently on order (Stalmokaitė and Hassler, 2020) This sluggish uptake is largely due to the long operational lifespan of vessels, typically 25-30 years, which creates significant inertia in the industry. Upgrading existing ships or investing in new builds with advanced technologies requires substantial capital, making fleet-wide changes extremely costly and time-consuming (Stalmokaitė et al., 2023). The capital-intensive nature of these investments leads to a lock-in effect, where shipping companies are hesitant to abandon existing technologies before fully deprecating their assets. This economic reality is compounded by the slow pace of regulatory change in the industry. For instance, despite recognizing the need to address greenhouse gas (GHG) emissions in the 1990s, the IMO did not adopt its first strategy for GHG reduction until 2018 (Stalmokaitė and Hassler, 2020). The combination of long vessel lifespans, high capital costs, and a slow-moving regulatory environment creates a gap between the rapid evolution of new technologies and their implementation in the shipping industry.
5.5. A culture of secrecy
@@
The maritime industry grapples with a unique challenge regarding cybersecurity incident transparency and information sharing. In stark contrast to the aviation sector, which has cultivated a culture of open reporting and collaborative information exchange on safety and security matters, the maritime domain often shrouds cyber incidents in secrecy, primarily due to concerns in reputational damage (Nimmich and Goward, 2007). In the United States alone, at least eighteen federal agencies have responsibility for regulating various aspects of maritime transportation, with little to no formal methods of coordinating their efforts (for a Study of the Federal Role in the Marine Transportation System, 2004). This regulatory fragmentation contributes to the ”highly fragmented” and ”near chaotic” nature of the maritime domain (for a Study of the Federal Role in the Marine Transportation System, 2004). This lack of coordination potentially discourages transparent reporting, and this may hinder the industry’s collective ability to learn from past incidents and implement effective preventive measures against future cyber threats. Moreover, the opaque nature of maritime operations and regulations poses additional obstacles. For instance, it impedes the development of crucial technologies such as autonomous collision avoidance systems (van de Merwe et al., 2023), thereby stunting the overall advancement of maritime technology. Overcoming this ingrained secrecy to foster greater transparency and information sharing will require significant shifts in industry practices, regulations, and mindsets. Key stakeholders must recognize that the benefits of shared knowledge and collaborative security efforts far outweigh the perceived risks of disclosure. True security lies not in secrecy, but in transparency.
5.6. Limitations of technological solutions
@@
As previously discussed, while technologies such as blockchain, machine learning, and artificial intelligence show promise in addressing maritime cybersecurity challenges, they are not panaceas. Each technology comes with its own set of limitations, including scalability issues, susceptibility to new forms of attacks, and integration difficulties with existing systems. For example, the global scale and complexity of maritime supply chains strain blockchain’s scalability and processing speed, potentially slowing down operations in an industry where time is critical (Li et al., 2020). It is essential for the maritime industry to approach these technologies carefully, recognizing both their potential benefits and inherent constraints. Additionally, rather than relying solely on a single advanced technology, a more effective strategy involves implementing a multi-layered technological approach. This may entail combining various technologies to establish a stronger defense. For instance, utilizing AI for threat detection, blockchain for secure data sharing, and quantum-resistant cryptography for future-proofing communications. This layered approach can help mitigate the weaknesses of individual technologies. Overall, while advanced technologies hold promise for enhancing cybersecurity in the maritime sector, it is crucial to carefully evaluate their capabilities and limitations before implementation.
6. Future Directions
@@ 6.1. Maritime Cloud
The maritime cloud concept is currently in its formative stages, with ongoing projects and evolving implementations shaping its development. Park (Park, 2017) mentioned that the maritime cloud is closely tied to the development of e-Navigation, a strategic initiative spearheaded by the IMO to harmonize and enhance navigation systems globally. The integration of the maritime cloud with e-Navigation strategies could offer a opportunity to facilitate seamless information exchange across diverse maritime stakeholders. It has the potential to transcend the limitations of specific communication systems or channels, creating a unified platform that connects vessels, ports, coastal authorities and other maritime actors. Cloud technology offers numerous other advantages in the maritime domain, including enhanced data management (Cristea et al., 2017), improved maritime safety and security with distributed smart surveillance systems (Lee and Park, 2017), cost reduction (Ristov et al., 2014), system compatibility (Park, 2017), and real-time monitoring of vessel positions and status (Joszczuk-Januszewska, 2012). Despite these promising applications, further research is imperative across multiple domain within maritime cloud computing, such as security concerns in cloud implementation (Botunac et al., 2017).
6.2. Testbeds
@@
Maritime-specific cybersecurity testbeds can be used to simulate cyber attacks on ship systems, allowing for the development and testing of defense mechanisms without risking actual vessels or infrastructure. In response to the growing need for advanced testing and research capabilities in the maritime sector, several notable testbeds have been established worldwide. The eMaritime Integrated Reference Platform (eMIR) (Hahn, 2015), developed in Germany, combines both virtual and physical components for providing a robust environment to test maritime cyber-physical systems. In the United Kingdom, the University of Plymouth hosts the Cyber-SHIP Lab (Tam et al., 2019), a specialised facility for researchers to carry out cyber attack experiments. Meanwhile, in Singapore, the Singapore Polytechnic houses the Advanced Navigation Research Simulator (ANRS) (of Excellence in Maritime safety, 2021), which is designed to support maritime training and education. Existing testbeds rely heavily on simulation rather than real maritime hardware, as physical testbeds incorporating actual maritime equipment are costly to construct and maintain (Yamin et al., 2020)(Zhou, 2024). Each testbed architecture has its distinct advantages and limitations. Virtual testbeds enable testing of potentially dangerous scenarios without risking damage to physical equipment or endangering personnel (Longo et al., 2023b); however, they often struggle to accurately replicate wireless communications (Amro and Gkioulos, 2021), which are crucial in maritime settings. There is a need for a balance between virtual and physical components to achieve greater realism in maritime testbeds. While current testbeds are valuable, we believe they require continued investment and innovation to better support the evolving needs of the maritime industry.
7. Conclusion
@@ Appendix A APPENDIX
In this section, we have included supplementary information for our survey. Firstly, we present definitions for a list of discussed vessel components. We further analyzed their vulnerabilities, highlighting the possible attacks that can happen. Lastly, we have provided a list of all the acronyms used in this survey.
A.1. Definitions of Vessel Components
@@
AIS: The AIS is a Very High Frequency (VHF) radio broadcasting system that is used for traffic monitoring (Vessel Traffic Services), collision avoidance, search and rescue operations, aids to navigation, weather forecast (Australian Maritime Safety Authority) and accident investigation (Bothur et al., 2017)(Rivkin, 2023).
@@
VDR: VDR is an essential system that acts as a ”black box” to record key parametric data about a ship’s voyage, which can then be recovered and analyzed to investigate the causes of a maritime accident. VDR continuously stores data related to the status, command, and control of the ship from various sensors and systems, including date/time, ship’s position, speed, heading, bridge audio, radio communications, radar data, hull openings status, watertight doors status, etc (Piccinelli and Gubian, 2013).
RADAR: RADAR is an object detection system that emits radio waves and analyzes the reflected signals to detect the presence, direction, and range of objects in the surrounding environment (Awan and Al Ghamdi, 2019). It is indispensable for maintaining situational awareness and ensuring the safety of maritime operations, particularly in low visibility conditions or when navigating through congested waters.
ECDIS: ECDIS is a computer-based navigation system that serves as a digital alternative to traditional paper nautical charts, providing navigators with real-time, interactive, and comprehensive information about the ship’s position, course, and surrounding environment. It combines information from multiple navigational aids and sensors, including GNSS, Radar, and AIS (Rivkin, 2023). This information is superimposed on digital charts and empowers navigation officers to make informed decisions and maintain situational awareness.
@@
GMDSS: The GMDSS was developed by the International Maritime Organization (IMO) and became fully operational in 1999. It is designed to perform the main functions such as transmitting ship-to-shore distress alerts, coordinating search and rescue operations and providing general radiocommunications (Ilcev, 2020). The implementation of GMDSS has significantly improved maritime safety and rescue operations globally, providing a standardized system for distress communication and alerting across international waters.
A.2. Vulnerabilities of Vessel Components
@@
Vulnerability of AIS: It was developed in the 1990s with a primary focus on enhancing maritime safety and navigation, but its design did not thoroughly consider the potential risks posed by active cyber attacks or malicious actors seeking to exploit vulnerabilities in the system. Numerous known security flaws exist (Goudossis and Katsikas, 2019)(Kessler and Shepard, 2020):
@@
  • Since AIS does not have mechanisms to guarantee the integrity of sent data, it is impossible to confirm that the data sent by an AIS device is accurate and has not been tampered with. This weakness makes it possible for a vessel to falsely portray its attributes, including size or type.
  • AIS-broadcast communications lack encryption. Openly available AIS data may be found on a number of websites, including Marine Traffic, and can offer a wealth of information about ships and their travels, including specifics about the ship’s identity, cargo, and current position (Bothur et al., 2017)(Hall et al., 2015).
Vulnerability of GNSS: The main GNSS vulnerabilities that have been thoroughly examined are the radio frequency hazards associated with jamming and spoofing. Extreme space weather events have been identified as a significant vulnerability for GNSS, in addition to radio frequency concerns. In contrast to physical assaults, which are mostly theoretical because mature nation-states demand highly developed capabilities, extreme space weather phenomena represent a genuine and continuous threat to GNSS satellites and their communications. These occurrences are real threats, even though their frequency and intensity vary over the course of the 11-year solar cycle (Owens et al., 2021). There are several reasons why GNSS are susceptible (Drumhiller et al., 2017)(Schmidt et al., 2016):
@@
Vulnerability of VDR: VDR data could be contested in court and declared inadmissible as evidence if it turns out to be easily altered or manipulated. Legal proceedings pertaining to maritime accidents, insurance claims, and liability issues may be significantly impacted by this (Ashour, 2013). Consequently, VDR may be exposed to hazards associated with data availability, data integrity problems, and confidentiality breaches.
Vulnerability of RADAR: A number of subsystems, including antennas, transmitters, receivers, processors, and displays, make up the intricate marine RADAR systems (Cohen et al., 2019). There could be vulnerabilities in every component, which would increase the number of attack surfaces. However, given the specialized nature of marine radar systems and the challenges associated with conducting thorough security assessments in real-world scenarios, there might be fewer opportunities to thoroughly inspect the system for vulnerabilities. The task of simulating actual circumstances for security testing is hampered by the complicated operational settings, which include the sea environment, weather, and the requirement for uninterrupted functionality.
Vulnerability of ECDIS: The vulnerabilities in ECDIS are largely caused by a combination of technological factors and organisational issues as the maritime industry has been slow to adapt to the rapidly evolving cyber threat landscape. The main vulnerabilities identified in ECDIS are (Svilicic et al., 2019a)(Svilicic et al., 2019b):
@@
  • Insecure network configuration and services, such as directory traversal, unsafe HTTP methods allowed, and header injection vulnerabilities in the Apache web server (Dyryavyy, 2014).
  • Identical hardware/software used for primary and backup ECDIS: If one ECDIS workstation is compromised, the other is equally vulnerable to the same threat
@@
Vulnerability of VSAT: Many maritime operators appear to be unaware of the risks associated with transmitting sensitive data over satellite links without adequate security measures (Wu et al., 2018), and numerous maritime VSAT networks continue to use outdated and insecure protocols (Pavur et al., 2020). The author (Santamarta, 2018) observes a trend where historically isolated devices, such as VSAT systems, are now being designed with additional communication technologies like WiFi and Bluetooth. Although not explicitly identified as a cause of vulnerabilities, this trend implies that increased connectivity may expand the attack surface if proper security measures are not implemented.
@@
Vulnerability of GMDSS: The radio and satellite communications used by GMDSS can be intentionally jammed or unintentionally interfered with, which could disrupt its functionalities (Tam and Jones, 2018) GMDSS relies on various interconnected systems. A vulnerability in one system could potentially affect the entire GMDSS functionality. Systems like GPS, which GMDSS relies on for positioning, can be spoofed to provide false location data (Ilcev, 2020).
A.3. Acronyms
Table 6. Acronyms
Acronym Description
ABS American Bureau of Shipping
ADS-B Automatic Dependent Surveillance-Broadcast

### v3

@@ V-A Device heterogeneity
Maritime devices operate in a wide range of environments, ranging from ports and coastal areas and vast expanses of open seas. The diversity of maritime environments requires a corresponding diversity in device types. While this heterogeneity presents challenges for manufacturers and operators alike – including issues related to standardization and interoperability – it also ensures that appropriate technologies can be effectively deployed across various maritime scenarios. The maritime industry is facing an increasing demand for improved operational visibility at sea [155]. Enhanced operational visibility at sea contributes to overall safety by providing early detection of anomalies or hazards that could pose a threat to crew members or the environment. Maritime IoT addresses this requirement by making it possible to monitor vital indicators like fuel usage, machinery performance, and general vessel operations in real time. As a result, maritime stakeholders are able to make more informed decisions, optimize resource utilization, and improve overall efficiency and safety. However, the implementation of maritime IoT presents complexities. The system must accommodate a wide range of machine-type communication devices, from low-cost units with limited functionality to high-end devices offering advanced features [155]. Low-cost devices such as sensors and buoys often operate under power and energy constraints, posing challenges in selecting appropriate communication technologies and protocols that prioritize energy efficiency and longevity over high data rates or complex functionalities [155]. To address these challenges posed by device heterogeneity, edge computing and cloud-based solutions have been proposed [156]. While cloud-based solutions and edge computing offer valuable tools for managing device heterogeneity in maritime IoT, cloud solutions require reliable internet connections, which may not always be available in open seas where face connectivity issues [157]. In addition, implementing both cloud and edge solutions adds another layer of complexity to an already complex system of heterogeneous devices.
V-B A culture of secrecy
@@
The maritime industry grapples with a unique challenge regarding cybersecurity incident transparency and information sharing. In stark contrast to the aviation sector, which has cultivated a culture of open reporting and collaborative information exchange on safety and security matters, the maritime domain often shrouds cyber incidents in secrecy, primarily due to concerns in reputational damage [158]. In the United States alone, at least eighteen federal agencies have responsibility for regulating various aspects of maritime transportation, with little to no formal methods of coordinating their efforts [159]. This regulatory fragmentation contributes to the ”highly fragmented” and ”near chaotic” nature of the maritime domain [159]. This lack of coordination potentially discourages transparent reporting, and this may hinder the industry’s collective ability to learn from past incidents and implement effective preventive measures against future cyber threats. Moreover, the opaque nature of maritime operations and regulations poses additional obstacles. For instance, it impedes the development of crucial technologies such as autonomous collision avoidance systems [160], thereby stunting the overall advancement of maritime technology. Overcoming this ingrained secrecy to foster greater transparency and information sharing will require significant shifts in industry practices, regulations, and mindsets. Key stakeholders must recognize that the benefits of shared knowledge and collaborative security efforts far outweigh the perceived risks of disclosure. True security lies not in secrecy, but in transparency.
V-C Rapid technology evolution v.s. Slow-paced industry
The slow pace of technological adoption in the global shipping fleet is evident in the statistics: only 0.3% of operating vessels worldwide have implemented alternative energy solutions, with this figure rising slightly to 6.05% for ships currently on order [161] This sluggish uptake is largely due to the long operational lifespan of vessels, typically 25-30 years, which creates significant inertia in the industry. Upgrading existing ships or investing in new builds with advanced technologies requires substantial capital, making fleet-wide changes extremely costly and time-consuming [52]. The capital-intensive nature of these investments leads to a lock-in effect, where shipping companies are hesitant to abandon existing technologies before fully deprecating their assets. This economic reality is compounded by the slow pace of regulatory change in the industry. For instance, despite recognizing the need to address greenhouse gas (GHG) emissions in the 1990s, the IMO did not adopt its first strategy for GHG reduction until 2018 [161]. The combination of long vessel lifespans, high capital costs, and a slow-moving regulatory environment creates a gap between the rapid evolution of new technologies and their implementation in the shipping industry.
V-D Limitations of technological solutions
@@
As previously discussed, while technologies such as blockchain, machine learning, and artificial intelligence show promise in addressing maritime cybersecurity challenges, they are not panaceas. Each technology comes with its own set of limitations, including scalability issues, susceptibility to new forms of attacks, and integration difficulties with existing systems. For example, the global scale and complexity of maritime supply chains strain blockchain’s scalability and processing speed, potentially slowing down operations in an industry where time is critical [162]. It is essential for the maritime industry to approach these technologies carefully, recognizing both their potential benefits and inherent constraints. Additionally, rather than relying solely on a single advanced technology, a more effective strategy involves implementing a multi-layered technological approach. This may entail combining various technologies to establish a stronger defense. For instance, utilizing AI for threat detection, blockchain for secure data sharing, and quantum-resistant cryptography for future-proofing communications. This layered approach can help mitigate the weaknesses of individual technologies. Overall, while advanced technologies hold promise for enhancing cybersecurity in the maritime sector, it is crucial to carefully evaluate their capabilities and limitations before implementation.
VI Future Directions
@@ VI-A Maritime Cloud
The maritime cloud concept is currently in its formative stages, with ongoing projects and evolving implementations shaping its development. Park [163] mentioned that the maritime cloud is closely tied to the development of e-Navigation, a strategic initiative spearheaded by the IMO to harmonize and enhance navigation systems globally. The integration of the maritime cloud with e-Navigation strategies could offer a opportunity to facilitate seamless information exchange across diverse maritime stakeholders. It has the potential to transcend the limitations of specific communication systems or channels, creating a unified platform that connects vessels, ports, coastal authorities and other maritime actors. Cloud technology offers numerous other advantages in the maritime domain, including enhanced data management [164], improved maritime safety and security with distributed smart surveillance systems [165], cost reduction [166], system compatibility [163], and real-time monitoring of vessel positions and status [167]. Despite these promising applications, further research is imperative across multiple domain within maritime cloud computing, such as security concerns in cloud implementation [168].
VI-B Testbeds
@@
Maritime-specific cybersecurity testbeds can be used to simulate cyber attacks on ship systems, allowing for the development and testing of defense mechanisms without risking actual vessels or infrastructure. In response to the growing need for advanced testing and research capabilities in the maritime sector, several notable testbeds have been established worldwide. The eMaritime Integrated Reference Platform (eMIR) [169], developed in Germany, combines both virtual and physical components for providing a robust environment to test maritime cyber-physical systems. In the United Kingdom, the University of Plymouth hosts the Cyber-SHIP Lab [170], a specialised facility for researchers to carry out cyber attack experiments. Meanwhile, in Singapore, the Singapore Polytechnic houses the Advanced Navigation Research Simulator (ANRS) [171], which is designed to support maritime training and education. Existing testbeds rely heavily on simulation rather than real maritime hardware, as physical testbeds incorporating actual maritime equipment are costly to construct and maintain [172][173]. Each testbed architecture has its distinct advantages and limitations. Virtual testbeds enable testing of potentially dangerous scenarios without risking damage to physical equipment or endangering personnel [174]; however, they often struggle to accurately replicate wireless communications [175], which are crucial in maritime settings. There is a need for a balance between virtual and physical components to achieve greater realism in maritime testbeds. While current testbeds are valuable, we believe they require continued investment and innovation to better support the evolving needs of the maritime industry.
VII Conclusion
@@ Appendix A APPENDIX
In this section, we have included supplementary information for our survey. Firstly, we present definitions for a list of discussed vessel components. We further analyzed their vulnerabilities, highlighting the possible attacks that can happen.
A-A Definitions of Vessel Components
@@
AIS: The AIS is a Very High Frequency (VHF) radio broadcasting system that is used for traffic monitoring (Vessel Traffic Services), collision avoidance, search and rescue operations, aids to navigation, weather forecast (Australian Maritime Safety Authority) and accident investigation [176][177].
@@
VDR: VDR is an essential system that acts as a ”black box” to record key parametric data about a ship’s voyage, which can then be recovered and analyzed to investigate the causes of a maritime accident. VDR continuously stores data related to the status, command, and control of the ship from various sensors and systems, including date/time, ship’s position, speed, heading, bridge audio, radio communications, radar data, hull openings status, watertight doors status, etc [178].
RADAR: RADAR is an object detection system that emits radio waves and analyzes the reflected signals to detect the presence, direction, and range of objects in the surrounding environment [179]. It is indispensable for maintaining situational awareness and ensuring the safety of maritime operations, particularly in low visibility conditions or when navigating through congested waters.
ECDIS: ECDIS is a computer-based navigation system that serves as a digital alternative to traditional paper nautical charts, providing navigators with real-time, interactive, and comprehensive information about the ship’s position, course, and surrounding environment. It combines information from multiple navigational aids and sensors, including GNSS, Radar, and AIS [177]. This information is superimposed on digital charts and empowers navigation officers to make informed decisions and maintain situational awareness.
@@
GMDSS: The GMDSS was developed by the International Maritime Organization (IMO) and became fully operational in 1999. It is designed to perform the main functions such as transmitting ship-to-shore distress alerts, coordinating search and rescue operations and providing general radiocommunications [180]. The implementation of GMDSS has significantly improved maritime safety and rescue operations globally, providing a standardized system for distress communication and alerting across international waters.
A-B Vulnerabilities of Vessel Components
@@
Vulnerability of AIS: It was developed in the 1990s with a primary focus on enhancing maritime safety and navigation, but its design did not thoroughly consider the potential risks posed by active cyber attacks or malicious actors seeking to exploit vulnerabilities in the system. Numerous known security flaws exist [138][181]:
@@
  • Since AIS does not have mechanisms to guarantee the integrity of sent data, it is impossible to confirm that the data sent by an AIS device is accurate and has not been tampered with.
  • AIS-broadcast communications lack encryption. Openly available AIS data may be found on a number of websites, including Marine Traffic, and can offer a wealth of information about ships and their travels, including specifics about the ship’s identity, cargo, and current position [176][133].
Vulnerability of GNSS: The main GNSS vulnerabilities that have been thoroughly examined are the radio frequency hazards associated with jamming and spoofing. Extreme space weather events have been identified as a significant vulnerability for GNSS, in addition to radio frequency concerns. In contrast to physical assaults, which are mostly theoretical because mature nation-states demand highly developed capabilities, extreme space weather phenomena represent a genuine and continuous threat to GNSS satellites and their communications. These occurrences are real threats, even though their frequency and intensity vary over the course of the 11-year solar cycle [182]. There are several reasons why GNSS are susceptible [144][183]:
@@
Vulnerability of VDR: VDR data could be contested in court and declared inadmissible as evidence if it turns out to be easily altered or manipulated. Legal proceedings pertaining to maritime accidents, insurance claims, and liability issues may be significantly impacted by this [184]. Consequently, VDR may be exposed to hazards associated with data availability, data integrity problems, and confidentiality breaches.
Vulnerability of RADAR: A number of subsystems, including antennas, transmitters, receivers, processors, and displays, make up the intricate marine RADAR systems [185]. There could be vulnerabilities in every component, which would increase the number of attack surfaces. However, given the specialized nature of marine radar systems and the challenges associated with conducting thorough security assessments in real-world scenarios, there might be fewer opportunities to thoroughly inspect the system for vulnerabilities. The task of simulating actual circumstances for security testing is hampered by the complicated operational settings, which include the sea environment, weather, and the requirement for uninterrupted functionality.
Vulnerability of ECDIS: The vulnerabilities in ECDIS are largely caused by a combination of technological factors and organisational issues as the maritime industry has been slow to adapt to the rapidly evolving cyber threat landscape. The main vulnerabilities identified in ECDIS are [103][104]:
@@
  • Insecure network configuration and services, such as directory traversal, unsafe HTTP methods allowed, and header injection vulnerabilities in the Apache web server [108].
  • Identical hardware/software used for primary and backup ECDIS.
@@
Vulnerability of VSAT: Many maritime operators appear to be unaware of the risks associated with transmitting sensitive data over satellite links without adequate security measures [152], and numerous maritime VSAT networks continue to use outdated and insecure protocols [99]. The author [186] observes a trend where historically isolated devices, such as VSAT systems, are now being designed with additional communication technologies like WiFi and Bluetooth.
@@
Vulnerability of GMDSS: The radio and satellite communications used by GMDSS can be intentionally jammed or unintentionally interfered with, which could disrupt its functionalities [111] GMDSS relies on various interconnected systems. A vulnerability in one system could potentially affect the entire GMDSS functionality. Systems like GPS, which GMDSS relies on for positioning, can be spoofed to provide false location data [180].
A-C Acronyms
Lastly, we have provided a list of all the acronyms (See Table V) used in this survey.
TABLE V: Acronyms
Acronym Description Acronym Description
ABS American Bureau of Shipping IMO International Maritime Organisation
ADS-B Automatic Dependent Surveillance-Broadcast IoT Internet of Things